OpenStudy (anonymous):

Part A: Maria rented a coat at $285 for 3 days. If she rents the same coat for 6 days, she has to pay a total rent of $510. Write an equation in the standard form to represent the total rent (y) that Maria has to pay for renting the coat for x days.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

i would say y=95x but 95x6 doesn't =510 it equals 570 so idk

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

so (3, 285) and (6, 510) are two points that satisfy the required equation

ganeshie8 (ganeshie8):

knw how to write the equation of line given two points ?

OpenStudy (anonymous):

um, not really..

ganeshie8 (ganeshie8):

remember the slope formula ?

OpenStudy (anonymous):

Oh yes! lol hold on

OpenStudy (anonymous):

3 = x1 6 = x2 285 = y1 510 = y2

ganeshie8 (ganeshie8):

yes ! find the slope, m

OpenStudy (anonymous):

510 - 285 = 225 6 - 3 = 3

OpenStudy (anonymous):

225/3 = 75

ganeshie8 (ganeshie8):

yep ! slope of line, m = 75

ganeshie8 (ganeshie8):

so the equation of line will be : y = mx + b y = 75x + b

ganeshie8 (ganeshie8):

you need to find the value of "b" and plugin above ^^

ganeshie8 (ganeshie8):

any idea how to find the value of "b" ? :)

OpenStudy (anonymous):

Part B: Write the equation obtained in Part A using function notation.

OpenStudy (anonymous):

f(x) = 75x + b

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

oh and I found a question and it's actually not a question.. and the person who asked the question is sort of using profanity.. -_-

OpenStudy (anonymous):

http://openstudy.com/study#/updates/53d2c1bde4b0f2f7e5e8f4ac heres the link to it, it needs to be closed because they are just messing around

ganeshie8 (ganeshie8):

we're not done with part A, you need to find the value of "b"

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

but idk how to do that

ganeshie8 (ganeshie8):

our equation is : y = 75x + b and you know that the given two points satisfy this equation, so simply plugin ANY one point and solve b

ganeshie8 (ganeshie8):

may be plugin this point (3, 285)

ganeshie8 (ganeshie8):

plug x = 3, y = 285 into ur equation of line

OpenStudy (anonymous):

75 * 3 = 225 285 - 225 = 60 60 = b

ganeshie8 (ganeshie8):

Yep! the required equation is ?

OpenStudy (anonymous):

y = 75x + 60

ganeshie8 (ganeshie8):

yes ! but thats in slope intercept form, they want equation in standard form know how to rearrange it into standard form ?

ganeshie8 (ganeshie8):

standard form is this : \[Ax + By=C\]

OpenStudy (anonymous):

ok but what is A?

ganeshie8 (ganeshie8):

y = 75x + 60 can be rearranged to : 75x - y = -60

ganeshie8 (ganeshie8):

thats the standard form ^^

ganeshie8 (ganeshie8):

is that clear ?

OpenStudy (anonymous):

not really..

ganeshie8 (ganeshie8):

y = 75x + 60 we want to change it to "Ax + By = C" form so subtract 75x both sides, what do u get ?

ganeshie8 (ganeshie8):

y = 75x + 60 -75x -75x

ganeshie8 (ganeshie8):

you get : y - 75x = 60 right ?

OpenStudy (anonymous):

wait i get it now!, okay thanks! and um, what for part b?

OpenStudy (anonymous):

Part B: Write the equation obtained in Part A using function notation.

ganeshie8 (ganeshie8):

good :) you're right for part B, its just : f(x) = 75x + 60

OpenStudy (anonymous):

ah okay, but there is one last part.. i know how to do it but i cant explain how to do it Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals.

OpenStudy (anonymous):

Find the y-intercept and plot it on the graph. Use the slope to find a second point and plot it. Draw the line that connects the two points. Would this be good?

OpenStudy (anonymous):

Even if I am good at something, I can't really explain how to do it...

ganeshie8 (ganeshie8):

that looks good, but u need to talk more about specific value of y intercept (60) and the slope (75)

ganeshie8 (ganeshie8):

step1 : The y intercept of this function is 60 which can be plotted on y axis : (0. 60) step2 : using slope (75) the second point can be plotted by going right 1 unit and up 75 units : (1, 135) step3 : draw a line through the above two points

OpenStudy (anonymous):

Okay, thank you so much, you helping be understand this alot better! :)